# Ruminations Upon the Modeling of X-ray Foregrounds, Backgrounds and Faint Sources

### v1

@@ 3.4 Out-of time events
3.5 Spectral modeling and deprojection of the ICM
Our strategy for spectrally modeling the cluster follows that of Mantz et al. (2014, 2016a). Given spectra extracted in a series of circular annuli covering the cluster (see Section 4.3), the ICM is modeled as a series of concentric, isothermal, spherical shells (each represented by an absorbed thermal emitter), with inner and outer radii corresponding to those of the annuli. In practice, this procedure allows the densities in each shell to be constrained individually, while temperatures and metallicities must be linked between groups of adjacent shells.
@@ 3.5 Spectral modeling and deprojection of the ICM
The geometric projection of the 3D model onto 2D annuli in the plane of the sky (Kriss et al., 1983) is handled by the projct model in xspec for Chandra data. For XMM, the additional mixing due to the PSF cannot be neglected, and so we implement the model more explicitly as a linear combination of thermally emitting components. Following Mantz et al. (2020), we employ the PSF model of Read et al. (2011) and compute the PSF mixing between annuli based on a $\beta$-model fit to the surface brightness distribution of the ICM from the Chandra data (see Section 4.1). In both cases, the projected emission from 3D radii larger than the extent of the final cluster annulus is accounted for post-facto by assuming a locally $\beta$-model profile (essentially equivalent to a power law at large radii). This final step often has a tens-of-percent impact on the density of the largest-radius shell modeled, but negligible effect at smaller radii (see Mantz et al. 2014).
4 Analysis of a given field
The first step in analyzing the data for a given cluster and the surrounding field is to obtain a baseline SFG model from offset ROSAT data, following Section 3.1. We then proceed with imaging and spectral analysis of the Chandra and/or XMM data as described below.
4.1 Determining the unresolved CXB component
As described in Section 3.2, a preliminary image model is needed in order to compute the residual CXB normalization as a function of position, after masking point-like sources identified in Chandra data. We use images in the 0.6–7.0 keV band and 2.25 keV exposure maps, matching those employed for source-finding, but binned by a factor of 4 for the sake of efficiency. Multiple observations are fit in parallel, using the correct Poisson likelihood for the number of counts in each pixel and masking point sources detected in Section 3.2. The image model at this stage consists of the CCD-specific QPB contribution, a spatially uniform, vignetted component (comprising the SFG and CXB) and a beta model representing the ICM. We convert the SFG model from Section 3.1 to an expectation for the foreground surface brightness, and adopt this value minus 3 times its assigned uncertainty as the minimum for the combined SFG+CXB model, adopting a log-uniform prior over larger values. The best-fitting model is used as the “background” to determine the sensitivity limit for detecting CXB sources throughout the unmasked parts of the field (Section 3.2).
@@ 4.1 Determining the unresolved CXB component
4.2 Testing the foreground and background models
@@
4.3 Definition of cluster centers and annular regions
Having checked the appropriateness of the foreground and background models, and potentially revised the prior on the SFG surface brightness, we continue analysis of the Chandra images, specifically binned by a factor of 2 and in the 0.6–2.0 keV band, where the QPB is relatively small compared to the focused X-ray signal. We first fit a model consisting of the SFG (uniform and vignetted), CXB (non-uniform and vignetted), QPB (non-uniform and not vignetted) and a beta model describing the ICM. The non-ICM components of this fitted model are interpreted as background for the purpose of finding the center of the cluster using the Symmetry-Peakiness-Alignment (SPA) algorithm applied to unbinned images (Mantz et al. 2015; in previous work, the Chandra blank sky data served as a background estimate instead). For relaxed clusters, the result is sometimes adjusted at the arcsecond level to lie at the center of the cool core. Fixing this as the center of the cluster model(s), we next test whether including a second $\beta$ component to describe the ICM substantially improves the goodness of fit, and retain the double-$\beta$ model if so. From this final model fit, we then define a set of annuli, approximately logarithmically spaced, spanning from the cluster center to the radius where the cluster signal represents a $2\sigma$ excess over the background in the 0.6–2.0 keV band (Figure 7). Spectra extracted from these regions, along with cluster-free regions if available, are used in the analysis described in Section 4.4.
@@ 4.3 Definition of cluster centers and annular regions
4.4 Final spectral analysis
@@ 5 Examples and performance
5.1 Chandra+XMM: MACS J1423.8+2404
Our first test case for the methods described in this work is MACS J1423, a redshift 0.539 galaxy cluster discovered in the ROSAT All-Sky Survey and first cataloged in the MAssive Cluster Sample (MACS; Ebeling et al. 2007) that has been observed to similar (and useful) depth with both Chandra and XMM (Table 1). Figures 24 and 68 illustrate various intermediate stages of our analysis for this cluster; in particular, Figure 8 includes a subset of the cluster-free spectra used to verify the adequacy of the foreground and background models. Worth noting for this cluster is the relatively small angular extent of the X-ray emission and its sharp surface brightness peak, both of which increase the salience of the PSF modeling for XMM. In addition, MACS J1423 lies behind a region of relatively bright and spatially variable Galactic emission (Figure 2), such that the SFG is clearly visible in the Chandra data (at least in this relatively early ACIS-S data, which predates the significant reduction in soft response). Following Section 4.2, the overall normalization of the SFG model used in the Chandra analysis is not constrained by a ROSAT-based prior.
@@
Figure 11 shows the comparison of deprojected profiles using the methods described in this work with blank-sky background estimates. Note that, unlike the last case, the SFG is not unusually bright in this field. Our new analysis therefore uses the standard modeling priors based on ROSAT data, while the old-style analysis uses blank-sky estimates without an additional SFG component. The two sets of constraints agree well generally, though the full modeling approach finds a slightly higher metallicity and lower temperature in the cluster center, as well as a slightly higher and better constrained density at large radii. Given that the cluster center is substantially brighter than the average background (though, we note, significantly fainter than the center of MACS J1423), the differences in this region are presumably due to the sparsity of information in the corresponding blank-sky spectra and/or the binning required to use the modified Cash statistic (Section 1).
@@ Appendix A ACIS Calibration Correction
While performing a preliminary analysis of the Chandra data for SPT J2215-3537, one of the first applications of this work (Stueber et. al, in preparation), we noticed a systematic offset between temperatures measured from the observations associated with the 2 proposals targeting this cluster. [^14] Watson et al. (2025) have noted the same issue in ACIS observations of A2029 spanning 2000–2023, and suggest a correction, notionally similar to ours, based on those data. These differ both in the date of observation and in the use of ACIS-I (observed in Aug. 2020) vs ACIS-S (observations spread between Dec. 2021 and Aug. 2023). To investigate whether this was a reproducible effect and its potential causes, we turned to observations of A1795, which span the duration of the Chandra mission and a range of focal plane temperatures. Specifically, our analysis uses only “on-axis” observations, where the cluster center is at or near the aimpoint on either chip I3 or S3, as opposed to the “offset” pointings that are among the regular calibration observations made since 2010. We further exclude ObsIDs 494, the only warm-focal-plane, on-axis observation from early in the mission, and 21839, which is excessively flared. The observations are summarized in Table 5.
@@
We employed the full forward modeling of foregrounds and backgrounds, as described above, but have verified that our results in this section change negligibly if we instead follow the standard practice of using a blank sky background, normalized according to the high-energy count rate. For simplicity, our model of the cluster emission in this case is a simple absorbed thermal emission model (phabs*apec) without deprojection. The results shown in this section use caldb 4.12.2, the latest relevant update at this writing, although our original investigation used an earlier version; our results are qualitatively consistent across versions since 4.11.0.
The left column of Figure 15 shows constraints on the apec temperature, abundance and normalization parameters from each observation in the central circle and an intermediate annulus; temperature measurements for all regions are shown in Figure 17. We see relatively stable behavior prior to 2015, and thereafter a noticeable reduction in temperature and metallicity, and an increase in normalization (the latter more pronounced in the central circular region). The reversal of these trends since 2022 is a product of the gain update in caldb 4.12.2. The ACIS-I observations appear to show greater scatter in these late observations compared with ACIS-S, while neither displays an obvious residual trend with focal plane temperature (color coding in the figure). We note that these trends largely vanish when excluding the lowest energies from the fit rather than using the full 0.6–7.0 keV band; however, it is necessary to remove all the data at energies <3<3 keV to convincingly accomplish this.
@@
The right columns of Figures 15 and 17 show the measured apec parameters for different regions in A1795 when the correction defined by this procedure is applied; we see that the trends have largely been removed, even if (as expected) some residual scatter remains, particularly in the ACIS-I data. Significantly, this is the case not just in the cluster center where the correction is defined, but also for the annuli at larger radii, extending into the isothermal part of the temperature profile. The exception is the outermost annulus ($180^{\prime\prime}$$300^{\prime\prime}$), where the correction appears to be too strong, resulting in a slightly increasing trend of temperature with time. This suggests a possible relationship with the correction for contamination on the optical blocking filter, which is expected to spatially vary along with the temperature of the filter itself. In particular, this pattern is expected to be relatively simple within a few arcmin of the nominal aimpoints on ACIS-I and -S, and vary more quickly at larger radii on ACIS-I and within $\sim 1^{\prime}$ of the edges of ACIS-S adjacent and opposite to ACIS-I (Vikhlinin 2004; cf Figure 14). The offset calibration pointings could be used to investigate this further, though we have not done so at this time.
@@
where $\Delta T$ is the difference between the focal plane temperature and 119.554C in C. If the above formulae would yield $\alpha<0$ for dates prior to 2020, we set $\alpha=0$ and $A=1$ (i.e., no correction at early times; see Figure 16). The correction can either be applied using the plabs model with parameters $\alpha$ and $A$ in xspec, or equivalently by directly multiplying the effective area curve by the function $A(E/1\,\mathrm{keV})^{\alpha}$.
@@
Naturally, one should treat any extrapolation of the correction model into the future, particularly for ACIS-I, with caution. In addition, as noted above, there may be additional complexity at $\gtrsim 3^{\prime}$ distance from the aimpoint. Finally, note that the data presented here can only test the correction’s efficacy over a limited range in (redshift-corrected) source temperature, $kT/(1+z)\sim 4.5$–7.0 keV, although the good performance over this range provides reason for optimism. At this writing, the key observation of time dependence in the measured temperatures in the core of A1795 has been verified by the ACIS calibration team (albeit prior to the release of caldb 4.12.2; private communication). Ultimately, we expect this appendix to be obviated by a future update to the Chandra calibration.

### v2

@@ III.4 Out-of time events
III.5 Spectral modeling and deprojection of the ICM
Our strategy for spectrally modeling the cluster follows that of Mantz et al. (2014, 2016a). Given spectra extracted in a series of circular annuli covering the cluster (see Section IV.3), the ICM is modeled as a series of concentric, isothermal, spherical shells (each represented by an absorbed thermal emitter), with inner and outer radii corresponding to those of the annuli. In practice, this procedure allows the densities in each shell to be constrained individually, while temperatures and metallicities must be linked between groups of adjacent shells.
@@ III.5 Spectral modeling and deprojection of the ICM
The geometric projection of the 3D model onto 2D annuli in the plane of the sky (Kriss et al., 1983) is handled by the projct model in xspec for Chandra data. For XMM, the additional mixing due to the PSF cannot be neglected, and so we implement the model more explicitly as a linear combination of thermally emitting components. Following Mantz et al. (2020), we employ the PSF model of Read et al. (2011) and compute the PSF mixing between annuli based on a $\beta$-model fit to the surface brightness distribution of the ICM from the Chandra data (see Section IV.1). In both cases, the projected emission from 3D radii larger than the extent of the final cluster annulus is accounted for post-facto by assuming a locally $\beta$-model profile (essentially equivalent to a power law at large radii). This final step often has a tens-of-percent impact on the density of the largest-radius shell modeled, but negligible effect at smaller radii (see Mantz et al. 2014).
IV Analysis of a given field
The first step in analyzing the data for a given cluster and the surrounding field is to obtain a baseline SFG model from offset ROSAT data, following Section III.1. We then proceed with imaging and spectral analysis of the Chandra and/or XMM data as described below.
IV.1 Determining the unresolved CXB component
As described in Section III.2, a preliminary image model is needed in order to compute the residual CXB normalization as a function of position, after masking point-like sources identified in Chandra data. We use images in the 0.6–7.0 keV band and 2.25 keV exposure maps, matching those employed for source-finding, but binned by a factor of 4 for the sake of efficiency. Multiple observations are fit in parallel, using the correct Poisson likelihood for the number of counts in each pixel and masking point sources detected in Section III.2. The image model at this stage consists of the CCD-specific QPB contribution, a spatially uniform, vignetted component (comprising the SFG and CXB) and a beta model representing the ICM. We convert the SFG model from Section III.1 to an expectation for the foreground surface brightness, and adopt this value minus 3 times its assigned uncertainty as the minimum for the combined SFG+CXB model, adopting a log-uniform prior over larger values. The best-fitting model is used as the “background” to determine the sensitivity limit for detecting CXB sources throughout the unmasked parts of the field (Section III.2).
@@ IV.1 Determining the unresolved CXB component
IV.2 Testing the foreground and background models
@@
IV.3 Definition of cluster centers and annular regions
Having checked the appropriateness of the foreground and background models, and potentially revised the prior on the SFG surface brightness, we continue analysis of the Chandra images, specifically binned by a factor of 2 and in the 0.6–2.0 keV band, where the QPB is relatively small compared to the focused X-ray signal. We first fit a model consisting of the SFG (uniform and vignetted), CXB (non-uniform and vignetted), QPB (non-uniform and not vignetted) and a beta model describing the ICM. The non-ICM components of this fitted model are interpreted as background for the purpose of finding the center of the cluster using the Symmetry-Peakiness-Alignment (SPA) algorithm applied to unbinned images (Mantz et al. 2015; in previous work, the Chandra blank sky data served as a background estimate instead). For relaxed clusters, the result is sometimes adjusted at the arcsecond level to lie at the center of the cool core. Fixing this as the center of the cluster model(s), we next test whether including a second $\beta$ component to describe the ICM substantially improves the goodness of fit, and retain the double-$\beta$ model if so. From this final model fit, we then define a set of annuli, approximately logarithmically spaced, spanning from the cluster center to the radius where the cluster signal represents a $2\sigma$ excess over the background in the 0.6–2.0 keV band (Figure 7). Spectra extracted from these regions, along with cluster-free regions if available, are used in the analysis described in Section IV.4.
@@ IV.3 Definition of cluster centers and annular regions
IV.4 Final spectral analysis
@@ V Examples and performance
V.1 Chandra+XMM: MACS J1423.8+2404
Our first test case for the methods described in this work is MACS J1423, a redshift 0.539 galaxy cluster discovered in the ROSAT All-Sky Survey and first cataloged in the MAssive Cluster Sample (MACS; Ebeling et al. 2007) that has been observed to similar (and useful) depth with both Chandra and XMM (Table 1). Figures 24 and 68 illustrate various intermediate stages of our analysis for this cluster; in particular, Figure 8 includes a subset of the cluster-free spectra used to verify the adequacy of the foreground and background models. Worth noting for this cluster is the relatively small angular extent of the X-ray emission and its sharp surface brightness peak, both of which increase the salience of the PSF modeling for XMM. In addition, MACS J1423 lies behind a region of relatively bright and spatially variable Galactic emission (Figure 2), such that the SFG is clearly visible in the Chandra data (at least in this relatively early ACIS-S data, which predates the significant reduction in soft response). Following Section IV.2, the overall normalization of the SFG model used in the Chandra analysis is not constrained by a ROSAT-based prior.
@@
Figure 11 shows the comparison of deprojected profiles using the methods described in this work with blank-sky background estimates. Note that, unlike the last case, the SFG is not unusually bright in this field. Our new analysis therefore uses the standard modeling priors based on ROSAT data, while the old-style analysis uses blank-sky estimates without an additional SFG component. The two sets of constraints agree well generally, though the full modeling approach finds a slightly higher metallicity and lower temperature in the cluster center, as well as a slightly higher and better constrained density at large radii. Given that the cluster center is substantially brighter than the average background (though, we note, significantly fainter than the center of MACS J1423), the differences in this region are presumably due to the sparsity of information in the corresponding blank-sky spectra and/or the binning required to use the modified Cash statistic (Section I).
@@ VI Summary and Conclusion
Adopting generative models for the various X-ray backgrounds allows inferences to be drawn on a statistically firm and transparent footing. Rather than asking whether the background spectrum extracted from a given observation and region is appropriate, or how to best propagate the “error” from subtracting it, we can instead debate the more scientifically oriented questions of whether the model and its priors are physically defensible and adequate to explain the data. The approach alone is not a panacea, as the backgrounds still present a fundamental limitation on what we can learn from the data. In particular, once the source of interest becomes fainter than Poisson scatter in the background realization in a given region and/or energy range, no methodology will be able to detect or constrain it with high significance.
@@ Appendix A ACIS Calibration Correction
While performing a preliminary analysis of the Chandra data for SPT J2215-3537, one of the first applications of this work (Stueber et al., 2026), we noticed a systematic offset between temperatures measured from the observations associated with the 2 proposals targeting this cluster. [^14] Watson et al. (2026) have noted the same issue in ACIS observations of A2029 spanning 2000–2023, and suggest a correction, notionally similar to ours, based on those data. These differ both in the date of observation and in the use of ACIS-I (observed in Aug. 2020) vs ACIS-S (observations spread between Dec. 2021 and Aug. 2023). To investigate whether this was a reproducible effect and its potential causes, we turned to observations of A1795, which span the duration of the Chandra mission and a range of focal plane temperatures. Specifically, our analysis uses only “on-axis” observations, where the cluster center is at or near the aimpoint on either chip I3 or S3, as opposed to the “offset” pointings that are among the regular calibration observations made since 2010. We further exclude ObsIDs 494, the only warm-focal-plane, on-axis observation from early in the mission, and 21839, which is excessively flared. The observations are summarized in Table 5.
@@
Figure 16: Temperatures measured for A1795, presented as in Figure 15, for the central region (left) and an annulus at intermediate radii (right). Energies <3<3 keV are ignored in these fits, at the expense of precision; however, no time-dependent correction appears necessary.
We employed the full forward modeling of foregrounds and backgrounds, as described above, but have verified that our results in this section change negligibly if we instead follow the standard practice of using a blank sky background, normalized according to the high-energy count rate. For simplicity, our model of the cluster emission in this case is a simple absorbed thermal emission model (phabs*apec) without deprojection. The results shown in this section use caldb 4.12.2 and ciao 4.17, the latest versions at this writing, although our original investigation used an earlier version; our results are qualitatively consistent across versions since 4.11.0.151515While this paper was in revision, ciao 4.18 and caldb 4.12.3 were released. This update provides a clear improvement over the uncorrected results shown in this section, although smaller time-dependent trends and the significant scatter in ACIS-I measurements remain. Table 6 provides our best correction for this newer calibration version as well.
The left column of Figure 15 shows constraints on the apec temperature, abundance and normalization parameters from each observation in the central circle and an intermediate annulus; temperature measurements for all regions are shown in Figure 18. We see relatively stable behavior prior to 2015, and thereafter a noticeable reduction in temperature and metallicity, and an increase in normalization (the latter more pronounced in the central circular region). The reversal of these trends since 2022 is a product of the gain update in caldb 4.12.2, and is largely corrected in 4.12.3 (see Footnote 20). The ACIS-I observations appear to show greater scatter in these late observations compared with ACIS-S, while neither displays an obvious residual trend with focal plane temperature (color coding in the figure). We note that these trends largely vanish when excluding the lowest energies from the fit rather than using the full 0.6–7.0 keV band; however, it is necessary to remove all the data at energies <3<3 keV to convincingly accomplish this (Figure 16).
@@
The right columns of Figures 15 and 18 show the measured apec parameters for different regions in A1795 when the correction defined by this procedure is applied; we see that the trends have largely been removed, even if (as expected) some residual scatter remains, particularly in the ACIS-I data. Significantly, this is the case not just in the cluster center where the correction is defined, but also for the annuli at larger radii, extending into the isothermal part of the temperature profile. The exception is the outermost annulus ($180^{\prime\prime}$$300^{\prime\prime}$), where the correction appears to be too strong, resulting in a slightly increasing trend of temperature with time. This suggests a possible relationship with the correction for contamination on the optical blocking filter, which is expected to spatially vary along with the temperature of the filter itself. In particular, this pattern is expected to be relatively simple within a few arcmin of the nominal aimpoints on ACIS-I and -S, and vary more quickly at larger radii on ACIS-I and within $\sim 1^{\prime}$ of the edges of ACIS-S adjacent and opposite to ACIS-I (Vikhlinin 2004; cf Figure 14). The offset calibration pointings could be used to investigate this further, though we have not done so at this time.
@@
where $\Delta T$ is the difference between the focal plane temperature and $119.554^{\circ}$C in C. If the above formulae would yield $\alpha<0$ for dates prior to 2020, we set $\alpha=0$ and $A=1$ (i.e., no correction at early times; see Figure 17). The correction can either be applied using the plabs model with parameters $\alpha$ and $A$ in xspec, or equivalently by directly multiplying the effective area curve by the function $A(E/1\,\mathrm{keV})^{\alpha}$.
@@
4.12.3 4.751 2.602 0.0 1.142516 7.3920-7.3920 3.042   1.6660 0.0 0.9891564 2.1841-2.1841
Naturally, one should treat any extrapolation of the correction model into the future, particularly for ACIS-I, with caution. In addition, as noted above, the correction as defined here may not be appropriate at radii $\gtrsim 3^{\prime}$ distance from the aimpoint. Finally, note that the data presented here can only test the correction’s efficacy over a limited range in (redshift-corrected) source temperature, $kT/(1+z)\sim 4.5$–7.0 keV, although the good performance over this range provides reason for optimism. At this writing, the key observation of time dependence in the measured temperatures in the core of A1795 has been verified by the ACIS calibration team (albeit prior to the release of caldb 4.12.2; private communication). Ultimately, we expect this appendix to be obviated by a future update to the Chandra calibration.